5) N2(g) + O2(g) 2 NO(g) ΔΗ 180.7 kJ ; AS°-24.7 J/K 25°C (a)Calculate the standard free-energy change for the reaction at 25 C. (b) Calculate AG at 500 C using the value calculated from part a. 6) (Hess Law) AG +1102 kJ/mole What is the change in free energy when 18 moles of chlorine gas are rea to make phosphorus trichloride cted with phosphorus
Answer #1

\Delta G^{0} = \Delta H^{0} - T\Delta S^{0}

a) AG 180.7KJ- 298K 0.0247KJ/K 173.34K.J

b) \frac{\delta G}{\delta T} = -S

OR, \Delta G_{2} = \Delta G_{1} - S (T2 -T1)

OR, \Delta G_{2} = 173.34 - 0.0247 *475 = 161.60 KJ

For THE REVERSE REACTION THE VALUE OF \Delta G ^{0} = -1102 KJ/mol

hence, for 18 mol it will be = -1102*18 = -19836 kj